# Cross-Camera Distracted Driver Classification through Feature Disentanglement and Contrastive Learning

### v1

@@ IV-B2 Backbones
We select three efficient backbone models to assess the impact of varying parameter counts on our method including GhostNet-v1.0 [32], MobileNetV3-S [35], and SqueezeNet-v1.1 [36].
@@ IV-B3 Implementation details
Our experiments are performed on a computer featuring an Intel i7-8700 CPU @3.20GHz and an NVIDIA GeForce GTX 1070 Ti. The operating system is Ubuntu 22.04, and the framework is PyTorch 2.0.1. We utilize the SGD optimizer with a momentum of 0.9 and a weight decay equal to $5\times 10^{-4}$. The batch size is set to 64. The GhostNet-v1.0 backbone is pretrained on ImageNet. The view queries are randomly initialized by sampling from a normal distribution with mean equal to 0 and variance 1.
@@ IV-C1 Ablation study
In this section we present the results detailing the contribution of the key components within the proposed model. Specifically, we compare three different efficient backbones and we assess the contribution of contrastive learning, the influence of the disentanglement module, and finally, the effect of their combined integration.
@@
Table II reports the comparison of the three model architectures in terms of number of parameters and number of floating point operations per second (FLOPS). The GhostNet-v1.0 model has 5.2 million parameters and 0.16 GFlops, which indicates a moderate computational complexity with a relatively higher number of parameters compared to the other models. On the other hand, the MobileNetV3-S model, demonstrates a significantly lower number of parameters at 2.5 million and the lowest computational complexity with 0.06 GFlops. Finally, the SqueezeNet-v1.1 model features the smallest number of parameters at 1.2 million but has the highest computational complexity among the three, with 0.35 GFlops.
@@
The results reported in Table III reveal notable performance differences among the considered backbones in the leave-one-camera-out experiments. GhostNet consistently demonstrates superior performance across most daytime and nighttime scenarios. For instance, it achieves the highest Top-1 accuracy of 62.51% and Top-5 accuracy of 93.75% when evaluating D2. This suggests that GhostNet is particularly effective in handling variations across different camera configurations. MobileNetV3 also performs well, especially during nighttime acquisitions, where it shows competitive results in some scenarios. For example, it achieves a Top-1 accuracy of 33.18% for N1 and 47.75% for N2, indicating its robustness in lower light conditions. However, its overall performance remains slightly lower than GhostNet’s across most evaluations. SqueezeNet, on the other hand, consistently underperforms compared to GhostNet and MobileNetV3. It struggles particularly in more challenging scenarios such as D4 and N4, where it records a notably low Top-1 accuracy of 7.91% and 4.22%, respectively. This highlights potential limitations of SqueezeNet in handling diverse and complex data conditions.
@@
TABLE IV: Ablation experiments about feature disentanglement and contrastive learning, using Top-1 and Top-5 (%) accuracy for leave-one-camera-out experiments.
@@ IV-C2 Comparison with other methods
In this section, we compare the results achieved by our DBMNet with those of four recent methods, namely Duan et al. [27], GhostNet [4], Hu et al. [5], and OLCMNet [16]. It is worth noting that Duan et al. trained a GhostNet-v1.0 by exploiting their proposed S-Softmax classifier for improving model robustness. GhostNet is the same architecture used in our method as the backbone, and unlike our method, it is trained exclusively through cross-entropy loss optimization for driver action classification. Hu et al. [5] trained a ResNet-18 using a compound loss, which includes the proposed Clustering Supervised Contrastive Loss (C-SCL) for improving action representations, along with the cross-entropy loss used for the classification head. For a fair comparison with our DBMNet, we replaced the ResNet18 architecture with our GhostNet-v1.0. Finally, OLCMNet is a lightweight network with an OLCM block to improve model effectiveness.
Table V reports the results for each round of our leave-one-camera-out protocol. According to the results, several considerations can be made. First, the worst result obtained by our model is for the D4 camera, Top-1 accuracy 11.29%, compared with the best result obtained for D2 of 62.51% Top-1 accuracy. This result was expected since the framing of the driver is very different from the other three cameras (see Figure 2 for a better understanding). Second, the cameras are ranked in descending order based on their Top-1 performance as follows: D2, D3, D1, and D4. This ranking serves as a valuable indicator of the viewpoints offering superior driver visibility, thereby guiding the optimal placement of cameras. Third, our DBMNet demonstrates superior performance compared to the other two methods with an average improvement of 9% across all four cameras. In particular, the most significant disparity is observed against GhostNet, where DBMNet achieves a remarkable 17.20% increase in Top-1 accuracy and a 16.44% increase in Top-5 accuracy on camera D3. Still, on camera D3, DBMNet exhibits the largest performance improvement over Hu et al. , with an improvement of 7.32% and 5.87% in Top-1 and Top-5 accuracy, respectively.
TABLE V: Top-1 and Top-5 (%) accuracy for our leave-one-camera-out experiments. D indicates acquisitions during the daytime. The number following the D indicates the camera ID. The names on the left side of the arrow refer to the training cameras, while the name on the right side of the arrow refers to the evaluated camera.
@@
Figure 3 shows the confusion matrices for each cross-camera configuration. For the daytime subset we can the following observations. Confirming that the best performance is obtained for D2 and D3, there is a prominent diagonal in the respective confusion matrices. The confusion matrix for D1 shows a high bias in the class “Hair / Makeup” (#7). Many of the test samples are wrongly categorized with this class. An analysis of the distribution of the samples per category, however, excludes the presence of an imbalance in sample cardinality that could favor any specific category. Therefore, the negative results can only be attributed to the poor generalization ability of the model on this viewpoint. In the case of D4, “Looking to the left” (#8) and “Reaching behind” (#19) consistently emerge as the most frequently predicted categories for the test samples, even when predictions are incorrect. With an accuracy rate of 53.00%, the class “Drinking / Eating (right)” stands as the second most accurately classified category for this camera. In general, for all daytime cameras, the “Looking Left” (#8) class is recognized with the highest rate. Confusion matrices for the nighttime subset reflect the poor performance, with only N1 and N2 presenting a discernible diagonal. The highest performance for N1 is observed in the “Reaching behind” (#19) class with 77% accuracy and the “Hair / Makeup” (#7) class with 71% accuracy. For N2, several classes achieve accuracy higher than or equal to 70%, namely “Talk with cellphone (left)” (#3), “Talk with cellphone (right)” (#4), “Looking Left” (#8), “Looking up” (#10), and “Operating GPS” (#18). In the case of N3, the model shows a high concentration of false positives for the “Normal driving” (#0) and “Reaching behind” (#19) classes. For N4, “Talk with cellphone (right)” (#4) and “Hair / Makeup” (#7) classes are recognized with accuracyies of 44% and 54%, respectively. The lower performance for N4 might be due to its significant difference from the other views, making it challenging to recognize actions accurately, with only these two actions being relatively easier to identify.
@@ IV-C3 Cross-dataset results
Since the camera configuration for AUCDD-V1, EZZ2021, and SFD is very similar to that of camera D4 in the 100-Driver dataset (see Figure 4), we perform inference on these three datasets using the model trained on images acquired from cameras D1, D2, and D3. Thus, we can state that the current setup constitutes a cross-camera and cross-dataset experiment. The experimental results shown in Table VI highlight the performance of different models on the D4 test set of 100-Driver and the AUCDD-V1, EZZ2021, and SFD datasets. It is evident from the results that DBMNet, our proposed model, outperforms the other models across most metrics. Specifically, DBMNet achieves the highest Top-1 accuracy on the D4 test set of 100-Driver (11.29%), AUCDD-V1 (7.71%), EZZ2021 (10.51%), and SFD (9.86%). Similarly, for Top-5 accuracy, DBMNet also shows superior performance on the D4 test set of 100-Driver (39.31%) and SFD (38.62%). In contrast, GhostNet, Hu et al. , and Duan et al. models exhibit lower Top-1 and Top-5 accuracies in most cases. For instance, GhostNet shows a Top-1 accuracy of 8.85% on the D4 test set of 100-Driver, which is lower than that of DBMNet. Similarly, the Top-5 accuracy of Hu et al. on EZZ2021 is only 21.53%, significantly less than the 34.09% achieved by DBMNet. Duan et al. ’s model performs comparatively better than GhostNet and Hu et al. , but still falls short of DBMNet, particularly in the SFD dataset where DBMNet achieves a Top-1 accuracy of 9.86% and a Top-5 accuracy of 38.62%. These results indicate that DBMNet’s training on the combination of D1, D2, and D3 cameras of the 100-Driver dataset provides a more robust model for cross-camera and cross-dataset applications, demonstrating improved generalization and higher accuracy across diverse datasets.
TABLE VI: Top-1 and Top-5 accuracy (%) on the D4 test set of 100-Driver, and the AUCDD-V1, EZZ2021 and SFD datasets for the model trained on the combination of D1, D2 and D3 cameras of the 100-Driver dataset.
@@ IV-C4 Analysis
In this section, we conduct an analysis of features to support our claim of disentangled features. To this end, we qualitatively evaluate how the features before and after disentanglement module separate in the embedding space and quantitatively estimate the capacity of discriminating viewpoints before and after feature disentanglement. Practically, we extract feature vectors before and after the View disentanglement module, namely $\mathbf{f}$ and $\mathbf{\hat{f}}$, for both training and validation images. We chose validation images because the camera view categories overlap with those of the training set, unlike those of the test set. Figure 5 illustrates the features of the validation set for the image subset D1, D2, D3 using t-Distributed Stochastic Neighbor Embedding (t-SNE). The colors represent different driver action classes. As depicted, the $\mathbf{f}$-features separate into small, distinct clusters, even for the same action class, likely due to the actions being captured by different cameras. In contrast, the $\mathbf{\hat{f}}$-features form larger clusters that more effectively distinguish between different actions regardless of the camera used, thereby demonstrating the effectiveness of the feature disentanglement module. We then exploit a Minimum Distance Classifier (MDC) to classify camera viewpoints. Specifically, we train an MDC classifier on $\mathbf{f}$ feature vectors, referred as $\mathrm{MDC}_{\mathbf{f}}$, and another MDC classifier on $\mathbf{\hat{f}}$ feature vectors, labeled as $\mathrm{MDC}_{\mathbf{\hat{f}}}$. Finally, we classify $\mathbf{f}$ validation samples with $\mathrm{MDC}_{\mathbf{f}}$ and $\mathbf{\hat{f}}$ validation samples with $\mathrm{MDC}_{\mathbf{\hat{f}}}$. Results are reported in Table VII.
TABLE VII: Results of an MDC on the validation set for the different cross-camera setups.

### v2

@@ IV-B2 Backbones
We select three efficient backbone models – GhostNet-v1.0 [44], MobileNetV3-S [50], and SqueezeNet-v1.1 [51] – to evaluate the impact of varying parameter counts on our method. These models are widely recognized as lightweight and computationally efficient, making them well-suited for on-edge deployment in embedded automotive systems [52, 18].
@@ IV-B3 Implementation details
Our experiments are performed on a computer featuring an Intel i7-8700 CPU @3.20GHz and an NVIDIA GeForce GTX 1070 Ti. The operating system is Ubuntu 22.04, and the framework is PyTorch 2.0.1. We utilize the SGD optimizer with a momentum of 0.9 and a weight decay equal to $5\times 10^{-4}$. The batch size is set to 64. The GhostNet-v1.0 backbone is pretrained on ImageNet. The view queries are randomly initialized by sampling from a normal distribution with mean equal to 0 and variance 1. The two hyperparameters in Eq. 5, $\lambda_{\rm{ac}}$ and $\lambda_{\rm{vc}}$, are both empirically set to 1.
@@ IV-C1 Ablation study
In this section, we present the results highlighting the contribution of the key components of the proposed model. Specifically, we compare three different efficient backbones, assess the contribution of contrastive learning, evaluate the effect of hyperparameter optimization for balancing the two triplet losses, examine the influence of the disentanglement module, and finally, analyze the effect of integrating all component together.
@@
Table II reports the comparison of the three model architectures in terms of number of parameters and number of floating point operations per second (FLOPS). The GhostNet-v1.0 model has 3.9 million parameters and 0.16 GFLOPS, which indicates a moderate computational complexity with a relatively higher number of parameters compared to the other models. On the other hand, the MobileNetV3-S model, demonstrates a significantly lower number of parameters at 2.5 million and the lowest computational complexity with 0.06 GFLOPS. Finally, the SqueezeNet-v1.1 model features the smallest number of parameters at 1.2 million but has the highest computational complexity among the three, with 0.35 GFLOPS.
@@
The results reported in Table III reveal notable performance differences among the considered backbones in the leave-one-camera-out experiments. GhostNet consistently outperforms the others across most daytime and nighttime scenarios, achieving, for example, a Top-1 accuracy of 62.51% and Top-5 accuracy of 93.75% on D2. This strong performance suggests that GhostNet is particularly effective in handling variations across different camera configurations compared to the other architectures. This aligns with its design goals – GhostNet was specifically developed to overcome the limitations of earlier architectures, such as MobileNet and SqueezeNet, by increasing feature diversity through the Ghost module while maintaining computational costs low. MobileNetV3 also performs well – particularly under nighttime conditions, with Top-1 accuracy of 33.18% for N1 and 47.75% for N2 – indicating its robustness in lower light conditions. However, its overall performance remains slightly lower than GhostNet’s across most evaluations maybe due to its aggressive optimization for efficiency. SqueezeNet, although highly compact, struggles significantly in challenging scenarios such as D4 and N4 (Top-1 accuracy of 7.91% and 4.22%, respectively), likely due to its limited feature diversity resulting from aggressive parameter reduction.
@@
Triplet loss hyperparameter tuning
The balance between the two triplet losses – $\mathcal{L}_{\rm{ac}}$ and $\mathcal{L}_{\rm{vc}}$ – plays a critical role for optimizing model performance. In this experiment, we assess Top-1 accuracy as a function of the weighting parameters $\lambda_{\rm{ac}}$ and $\lambda_{\rm{vc}}$ as defined in Eq. 5, by training on D1, D2, and D3 and testing on D4. As shown in Figure 3, the highest accuracy of 11.29% is achieved when both losses are equally weighted with $\lambda_{\rm{ac}}=\lambda_{\rm{vc}}=1$. We can also observe that performance slightly degrades when the weights of the two losses are similar but lower than one. In contrast, performance significantly degrades when the balance between the two losses is skewed, with the lowest accuracy of 5.68% observed at $\lambda_{\rm{ac}}=1$ and $\lambda_{\rm{vc}}=0.5$.
Figure 3: Accuracy for the model trained on D1, D2, and D3 and tested on D4, as a function of the hyperparameters $\lambda_{\rm{ac}}$ and $\lambda_{\rm{vc}}$ defined in the Eq. 5. The red cross indicates the optimal configuration that yields the highest accuracy.
@@
TABLE IV: Ablation experiments about feature disentanglement and contrastive learning, using Top-1 and Top-5 (%) accuracy for leave-one-camera-out experiments. The results in the first row of this table match those reported for GhostNet [8] in Table VI This is because GhostNet [8] was trained using only cross-entropy loss, without the inclusion of feature disentanglement or contrastive learning.
@@ IV-C2 Comparison with other methods
In this section, we compare the results achieved by our DBMNet with those of four recent methods, namely DGCCL [36], Duan et al. [31], GhostNet [8], Hu et al. [9], and OLCMNet [20]. DGCCL is based on a ViT-B/16 architecture and is optimized using a composite loss function that combines center loss and ArcFace loss to enhance domain generalization. It is worth noting that Duan et al. trained a GhostNet-v1.0 by exploiting their proposed S-Softmax classifier for improving model robustness. GhostNet is the same architecture used in our method as the backbone, and unlike our method, it is trained exclusively through cross-entropy loss optimization for driver action classification. Hu et al. [9] trained a ResNet-18 using a compound loss, which includes the proposed Clustering Supervised Contrastive Loss (C-SCL) for improving action representations, along with the cross-entropy loss used for the classification head. Finally, OLCMNet is a lightweight network with an OLCM block to improve model effectiveness.
TABLE V: Comparison of the proposed DBMNet with literature methods in terms of number of parameters and computational complexity.
Params GFLOPS
DGCCL [36] 85.8M 17.58
Duan et al. [31] 4.3M 0.16
GhostNet [8] 3.9M 0.16
Hu et al. [9] 11.2M 1.83
OLCMNet [20] 2.8M 0.68
DBMNet (ours) 4.0M 0.16
Tables V and VI present a comparison between the proposed DBMNet and state-of-the-art methods in terms of model complexity and performance. Table V highlights that DBMNet achieves a favorable balance between accuracy and efficiency, with only 4.0M parameters and 0.16 GFLOPS – comparable to lightweight models like GhostNet (3.9M, 0.16 GFLOPS) and significantly more efficient than Hu et al. (11.2M, 1.83 GFLOPS) and especially DGCCL (85.8M, 17.58 GFLOPS). Performance-wise (Table VI), DBMNet consistently outperforms all baselines across the leave-one-camera-out protocol. The lowest Top-1 accuracy for DBMNet is observed on camera D4 (11.29%), and the highest on camera D2 (62.51%). This variation is expected due to the distinct of D4 compared to the other cameras (see Figure 2). Ranking the cameras by Top-1 accuracy yields the order: D2, D3, D1, and D4 – underscoring the influence of viewpoint on driver behavior recognition.
Compared to DGCCL, which is the most computationally expensive method by a wide margin, DBMNet achieves comparable or superior accuracy in most cases while being over 20 times more efficient in terms of FLOPS and more than 21 times smaller in parameter count. Specifically, DBMNet outperforms DGCCL in 5 out of 8 test scenarios, with notable gains on N1 (+14.86% Top-1), N3 (+2.30% Top-1), and N4 (+3.74% Top-1). DGCCL performs competitively on D1, D2, and D3, but its high complexity makes it less suitable for deployment in resource-constrained environments.
Overall, DBMNet delivers strong accuracy with minimal computational overhead, offering a practical and effective solution for cross-view driver state recognition.
TABLE VI: Top-1 and Top-5 (%) accuracy for our leave-one-camera-out experiments. D indicates acquisitions during the daytime. The number following the D indicates the camera ID. The names on the left side of the arrow refer to the training cameras, while the name on the right side of the arrow refers to the evaluated camera. In each column, the best and second best results are marked in boldface and underlined, respectively.
@@
Figure 4 shows the confusion matrices for each cross-camera configuration. For the daytime subset we can the following observations. Confirming that the best performance is obtained for D2 and D3, there is a prominent diagonal in the respective confusion matrices. The confusion matrix for D1 shows a high bias in the class “Hair / Makeup” (#7). Many of the test samples are wrongly categorized with this class. An analysis of the distribution of the samples per category, however, excludes the presence of an imbalance in sample cardinality that could favor any specific category. Therefore, the negative results can only be attributed to the poor generalization ability of the model on this viewpoint. In the case of D4, “Looking to the left” (#8) and “Reaching behind” (#19) consistently emerge as the most frequently predicted categories for the test samples, even when predictions are incorrect. With an accuracy rate of 53.00%, the class “Drinking / Eating (right)” stands as the second most accurately classified category for this camera. In general, for all daytime cameras, the “Looking Left” (#8) class is recognized with the highest rate. Confusion matrices for the nighttime subset reflect the poor performance, with only N1 and N2 presenting a discernible diagonal. The highest performance for N1 is observed in the “Reaching behind” (#19) class with 77% accuracy and the “Hair / Makeup” (#7) class with 71% accuracy. For N2, several classes achieve accuracy higher than or equal to 70%, namely “Talk with cellphone (left)” (#3), “Talk with cellphone (right)” (#4), “Looking Left” (#8), “Looking up” (#10), and “Operating GPS” (#18). In the case of N3, the model shows a high concentration of false positives for the “Normal driving” (#0) and “Reaching behind” (#19) classes. For N4, “Talk with cellphone (right)” (#4) and “Hair / Makeup” (#7) classes are recognized with accuracies of 44% and 54%, respectively. The lower performance for N4 might be due to its significant difference from the other views, making it challenging to recognize actions accurately, with only these two actions being relatively easier to identify.
@@ IV-C3 Cross-dataset results
Since the camera configuration for AUCDD-V1, EZZ2021, and SFD is very similar to that of camera D4 in the 100-Driver dataset (see Figure 5), we perform inference on these three datasets using the model trained on images acquired from cameras D1, D2, and D3. Thus, we can state that the current setup constitutes a cross-camera and cross-dataset experiment. The experimental results shown in Table VII highlight the performance of different models on the D4 test set of 100-Driver and the AUCDD-V1, EZZ2021, and SFD datasets. It is evident from the results that DBMNet, our proposed model, outperforms the other models in terms of Top-1 accuracy across three of the four datasets. Specifically, it achieves the highest Top-1 accuracy on the D4 test set of 100-Driver (11.29%), AUCDD-V1 (7.71%), and SFD (9.86%). On EZZ2021, it ranks second, with a Top-1 accuracy only 1.43% lower than the best-performing model, DGCCL. In terms of Top-5 accuracy, DBMNet also shows superior performance on the D4 test set of 100-Driver (39.31%) and SFD (38.62%), although it is outperformed by DGCCL on AUCDD-V1 and EZZ2021. In contrast, GhostNet, Hu et al., and Duan et al. models exhibit lower Top-1 and Top-5 accuracies in most cases. For instance, GhostNet shows a Top-1 accuracy of 8.85% on the D4 test set of 100-Driver, which is lower than that of DBMNet. Similarly, the Top-5 accuracy of Hu et al. on EZZ2021 is only 21.53%, significantly less than the 34.09% achieved by DBMNet. Duan et al.’s model performs comparatively better than GhostNet and Hu et al., but still falls short of DBMNet, particularly in the SFD dataset where DBMNet achieves a Top-1 accuracy of 9.86% and a Top-5 accuracy of 38.62%. These findings underscore the effectiveness of DBMNet’s training on the combination of D1, D2, and D3 cameras of the 100-Driver dataset provides a more robust model for cross-camera and cross-dataset applications, demonstrating improved generalization and higher accuracy across diverse datasets.
TABLE VII: Top-1 and Top-5 accuracy (%) on the D4 test set of 100-Driver, and the AUCDD-V1, EZZ2021 and SFD datasets for the model trained on the combination of D1, D2 and D3 cameras of the 100-Driver dataset. In each column, the best and second best results are marked in boldface and underlined, respectively.
@@ IV-C4 Analysis
In this section, we conduct an analysis of features to support our claim of disentangled features. To this end, we qualitatively evaluate how the features before and after disentanglement module separate in the embedding space and quantitatively estimate the capacity of discriminating viewpoints before and after feature disentanglement. Practically, we extract feature vectors before and after the View disentanglement module, namely $\mathbf{f}$ and $\mathbf{\hat{f}}$, for both training and validation images. We chose validation images because the camera view categories overlap with those of the training set, unlike those of the test set. Figure 6 illustrates the features of the validation set for the image subset D1, D2, D3 using t-Distributed Stochastic Neighbor Embedding (t-SNE). The colors represent different driver action classes. As depicted, the $\mathbf{f}$-features separate into small, distinct clusters, even for the same action class, likely due to the actions being captured by different cameras. In contrast, the $\mathbf{\hat{f}}$-features form larger clusters that more effectively distinguish between different actions regardless of the camera used, thereby demonstrating the effectiveness of the feature disentanglement module. We then exploit a Minimum Distance Classifier (MDC) to classify camera viewpoints. Specifically, we train an MDC classifier on $\mathbf{f}$ feature vectors, referred as $\mathrm{MDC}_{\mathbf{f}}$, and another MDC classifier on $\mathbf{\hat{f}}$ feature vectors, labeled as $\mathrm{MDC}_{\mathbf{\hat{f}}}$. Finally, we classify $\mathbf{f}$ validation samples with $\mathrm{MDC}_{\mathbf{f}}$ and $\mathbf{\hat{f}}$ validation samples with $\mathrm{MDC}_{\mathbf{\hat{f}}}$. Results are reported in Table VIII.
TABLE VIII: Results of an MDC on the validation set for the different cross-camera setups.
@@
IV-C5 Discussion
In this section, we summarize the results of both the cross-dataset and cross-camera experiments, incorporating an analysis from a computational efficiency perspective. The cumulative accuracy plot in Figure 7 provides a comprehensive overview of the generalization capabilities of the evaluated methods. Among them, our proposed DBMNet consistently demonstrates a superior balance between effectiveness and efficiency. In terms of effectiveness, DBMNet delivers an average Top-1 accuracy improvement of 7% compared to prior approaches. This performance gap is particularly evident when compared to methods that rely on efficient backbones, namely Duan et al., GhostNet, Hu et al., and OLCMNet. In comparison with DGCCL, DBMNet achieves comparable average accuracy, with a 1.36% improvement. A more detailed analysis reveals that DBMNet consistently ranks either first or second across the evaluated conditions, while DGCCL is significantly penalized in nighttime scenarios.
When combining these results with the computational efficiency data reported in Table V, the advantages of DBMNet become even more evident. The model comprises only 4.0M parameters and requires just 0.16 GFLOPS – a substantially more compact architecture than those of competing methods. For example, although DGCCL achieves similar accuracy, it does so at a much higher computational cost, requiring 85.8M parameters and 17.58 GFLOPS. This corresponds to over 20 times more parameters and more than 100 times the computational load compared to DBMNet. Even in comparison to relatively lightweight models such as GhostNet and Hu et al., DBMNet remains more efficient while still delivering superior performance. This combination of accuracy and efficiency makes DBMNet particularly well-suited for deployment in resource-constrained embedded systems, such as those used in automotive environments.
Figure 7: Boxplot summarizing the Top-1 accuracy achieved by the compared methods across all cross-camera and cross-dataset experiments.
IV-D Deployment on embedded device
The proposed DBMNet and the competing methods for distracted driver classification are deployed on the Coral Dev Board 4GB, a commercially available ARM-based Single-Board Computer (SBC) designed specifically for edge AI applications [54]. This development board integrates a Google Edge TPU coprocessor, a specialized hardware accelerator optimized for efficient execution of deep learning models directly on the device.
The Edge TPU is an 8-bit fixed-point coprocessor capable of performing high-speed inference on quantized TensorFlow Lite models with low power consumption. This makes it particularly suitable for real-time, on-device inference tasks, such as monitoring driver distraction. The Coral Dev Board therefore enables compact, energy-efficient, and responsive deployment of the proposed system in embedded automotive environments.
All methods are deployed following post-training quantization using full integer quantization. This optimization technique converts both the model weights and activations from 32-bit floating-point to 8-bit fixed-point representations. It requires calibration with a representative dataset to determine optimal quantization parameters for the activations. For this purpose, we use the validation set of the 100-Driver dataset in a leave-one-camera-out configuration. A quantized version of each model is generated for every configuration of daylight data. Due to the limited computational resources of the Coral Dev Board, the DGCCL model could not be successfully deployed and is therefore excluded from this evaluation.
Performance is evaluated across five metrics:
  • Average error represents the classification error obtained across the four folds of the leave-one-camera-out experiment;
  • Model size refers to the size of the quantized model file (including architecture and weights), reported in megabytes (MB);
  • Memory usage is the average memory consumption (in MB) during inference, computed over 100 inference passes;
  • Inference time is the average time (in seconds) required for a single forward pass, again averaged over 100 runs;
  • Power consumption is estimated by measuring the energy used across 100 inference passes, scaled by the average inference time, and reported in milliampere-hours (mAh).
The radar plot reported in Figure 8 highlights DBMNet (ours) as the best-performing model, achieving the lowest average error while maintaining a compact model size, low memory usage, fast inference time, and minimal power consumption – making it ideal for real-time, embedded applications. Compared to other methods, DBMNet offers the best trade-off between accuracy and efficiency. Duan et al. and GhostNet are as lightweight and power-efficient as DBMNet but suffers from high error rates. Hu et al. is the largest and most resource-intensive without offering a corresponding gain in accuracy. OLCMNet fall between these extremes but do not match DBMNet’s overall balance.
Figure 8: Radar plot comparing the proposed DBMNet with competing methods deployed on the Coral Dev Board across five key metrics: average error, model size, memory usage, inference time, and power consumption.